# Open Problems, and Future Trends

### v1

@@ III-B2 Storage Resource Allocation Issues
Unlike the stable resource scheduling in cloud centers, edge servers are less inclined to share storage space and data due to specific requirements such as workload, latency, and privacy[128]. In emergencies, the inconsistent availability information of storage resources on edge servers may lead to a high data loss rate. To address this issue, Vaquero et al.[129] propose a method to replicate data on all nodes. However, the above method will increase the overall cost of the system and may degrade performance. Recent research work also involves edge caching technology in Metaverse scenarios. Cai et al.[130] design an efficient control policy that integrates computing, caching, and communication (3C) resources and implements the policy by implementing a multi-pipeline traffic control and 3C resource scheduling mechanism. To further optimize the online delivery performance of data-intensive services, the authors also proposed a database placement strategy based on throughput maximization and two effective database replacement strategies. On the other hand, Huynh et al.[131] propose an innovative DT solution that combines MEC and URLLC to support Metaverse applications. From the perspective of the 3C integration model, this solution optimizes edge caching policies, task offloading strategies, and the allocation of computing and communication resources to address the latency optimization problem in the Metaverse enabled by DT.
@@ III-C Interoperability
As a complex system, the Industrial Metaverse platform integrates various sensors and control subsystems with different hardware architectures and control components. If there is no support for interoperability, user identities and industrial equipment data will be limited to specific VSPs, contradicting the original intention of an interconnected Metaverse[132]. Therefore, interoperability of subsystems within the platform and interoperability across platforms are important considerations. Currently, solving the interoperability issues of the Industrial Metaverse mainly starts from two aspects: identity and identification authentication, and interface and integration framework.
@@
Identity and identification define unique codes for users and production equipment, while binding the identification with attributes such as device status and human behavior[133]. This forms the foundation for achieving interoperability. These attributes can be provided by physical devices in factories or created by enterprise personnel themselves. In different platforms, the same identity and identification correspond to the same static attributes. Gadekallu et al.[134] suggest that BC can facilitate the exchange of data between different sub-Metaverses. They proposed using cross-chain protocols to exchange data on multiple BCs located in different virtual worlds. Therefore, Li et al.[133] propose a MetaOpera, a BC-based interoperability protocol for the Metaverse. It supports cross-chain technology and on-chain/off-chain techniques for the interaction between MetaOpera and decentralized platforms as well as top centralized platforms. Compared to authentication technologies based on Sidechains[135], this method achieves lower interaction latency. Additionally, Ghirmai et al.[136] propose using Self-Sovereign Identity (SSI) technology, which replaces traditional identity verification with distributed identity verification using zero-knowledge proofs. For example, two workers docked on a Metaverse platform can prove each other’s public keys according to the principles of SSI and encrypt their interaction records using pre-shared keys. Each worker holds a unique NFT as a unique identifier. However, the authors also pointed out that the issue of BC storage should not be ignored because controlling and managing data and identities may require different storage mechanisms. Additionally, their scenario only applies to static attributes in the Metaverse and does not involve behavioral objects.
Split learning, as a distributed architecture, involves multiple devices and servers in an industrial platform[137]. It can split the backbone network into multiple sub-networks, execute multiple sub-tasks in parallel on different devices or servers, and then integrate the results. Inspired by this, Rawal et al.[138] introduce a Split-Protocol that provides interoperability services. They assume that critical elements and functions are provided by a meta-server, and the system consists of ten identical meta-servers, each executing specific functions of the Metaverse. Role transitions can be achieved through the Split-Protocol. This protocol with role-switching capability can improve the availability of the Metaverse system. Furthermore, Hashash et al.[139] emphasize the coordination between DTs and Edge Metaverse. A Physical Twin (PT) system running in a large-scale sensing area is replicated as a Cyber Twin (CT) system on MEC servers. The area is divided into distributed subgrids and transmitted remotely to MEC servers. Their designed interactive algorithm utilizes optimal transport theory, considering both computational power and data synchronization issues.
In summary, each Industrial Metaverse platform has different configurations, and there is still a lack of standard development patterns. Interoperability and compatibility are guarantees for providing a better operating experience in the Industrial Metaverse. Cooperation between companies and industries is needed to increase the content and value of the Industrial Metaverse and establish an open collaborative ecosystem[140].
TABLE IV: Research on Existing issues in the Industrial Metaverse.
@@
Security & Confidentiality Sensor-side Industrial Data Source [89] A data misuse prevention mechanism called DAPter. DAPter can substantially raise the bar of the data abuse difficulty with little impact on QoS and overhead. 2021
@@
Inherent Security Vulnerabilities of Enabling Technologies [111] AcoPalm, a palm print recognition system. AcoPalm is resistant to replay and mimicry attacks, and can achieve 96.22% authentication accuracy. 2022
[112] Zeta protocol. The authors explain how it can be used with the available interaction methods offered by head-mounted displays. 2020
@@
[122] The design of the DL-based auction. Auction improves the revenue while satisfying the individual rationality and incentive compatibility constraints. 2022
@@ IV Standardization
Since the concept of the Metaverse was proposed, efforts have been made to promote its industrial deployment and ensure interoperability and compatibility between different Industrial Metaverse platforms, as well as to reduce the cost of industrial system development and maintenance. Standardization work in this regard has been ongoing.
The MPEG-V standardization, initiated in 2008, defined the exchange of data at the interface between the real world and the virtual world, with the aim of standardizing 4D sensory effects and the virtual world[143]. The standard consists of seven parts, among which 23005.1 provides an overview of the architecture and specifies the relevant information representation for achieving interoperability. 23005.2-3 covers syntax and semantic information, including sensory device functionality and sensor functionality. 23005.5-6 standardize data formats and data types, which are particularly important in industrial multimodal data scenarios[144]. Subsequently, the IEEE Standards Association launched the P2888 project (Interface Networking and Physical World) and established the Synchronization Metaverse as a supplement to the ISO/IEC 23005 series of standards. IEEE 2888 includes four parts: specifications for sensor interfaces in networking and physical worlds, standards for actuator interfaces in networking and physical worlds, orchestration of digital synchronization between networking and physical worlds, and architecture for VR disaster response systems. The first three parts aim to provide general technology for DT or the Metaverse, while the fourth standard aims to provide specific applications for the above three standards[145].
In June 2022, the Metaverse Standards Forum (MSF) was initiated and established by the Khronos Group, with participants including Meta, Microsoft, Epic, NVIDIA, Qualcomm, Sony, and other 37 technology giants engaged in the field of the Metaverse, covering industries such as chip manufacturers and game companies. The organization has now gathered over 2,400 industry vendors and standard organizations to collaborate across the industry on building an open Metaverse, interoperability testing, open-source tools, deployment guidance, and more. In July of the same year, ISO/IEC JTC 1/SC 29 MPEG WG2 Market Need Ad Hoc Group (AHG) launched Metaverse case collection and technical analysis to identify MPEG efficient coding and related system standard technical requirements that can support future Metaverse experiences. In January 2023, the IEC Standardization Evaluation Group 15 on Metaverse (SEG 15) was established. Its specific work includes studying the standardization roadmap for the Metaverse field in conjunction with current standardization achievements and development trends, establishing timely contact with TC/SC/Syc (including JTC1) as well as ISO and other relevant organizations, and providing work proposals for the International Electrotechnical Commission Standards Management Board (IECSMB). Recently, the IEEE P3812 series of standards is in the drafting phase, which defines the requirements for the Metaverse identity framework[146]. The provided identity framework is intended for use across different Metaverse systems to meet the requirements for interoperability/interaction of entities across multiple Metaverse systems. Additionally, the IEEE P2048.101 series of standards define the general technical framework, components, integration, and major business processes for AR systems applied to mobile devices[147].
In the future, as Industrial Metaverse gradually takes off, the accompanying standards regarding interoperability, security and privacy, openness and scalability, cross-industry collaboration, and other aspects will become increasingly refined, forming a comprehensive standard system.
@@ V-A Flexible Communication Network Architecture
The seamless internet connectivity of the Metaverse largely relies on reliable broadband services, with holographic communication requiring real-time operation, sufficient bandwidth, and throughput[148]. While low latency and high reliability are the primary advantages of 5G/6G services, consistent latency requires a flexible end-to-end architecture. This is because the Industrial Metaverse requires real-time communication of video, audio, and data. Currently, there is limited research on the network architecture and signal processing aspects of 6G in the context of the Metaverse. Relevant skills/mathematical tools: Wireless communication theory, signal processing, 5G/6G network architecture design.

### v2

@@ III-B2 Storage Resource Allocation Issues
Unlike the stable resource scheduling in cloud centers, edge servers are less inclined to share storage space and data due to specific requirements such as workload, latency, and privacy. In emergencies, the inconsistent availability information of storage resources on edge servers may lead to a high data loss rate. To address this issue, Vaquero et al.[151] propose a method to replicate data on all nodes. However, the above method will increase the overall cost of the system and may degrade performance. Recent research work also involves edge caching technology in Metaverse scenarios. Cai et al.[152] design an efficient control policy that integrates computing, caching, and communication (3C) resources and implements the policy by implementing a multi-pipeline traffic control and 3C resource scheduling mechanism. To further optimize the online delivery performance of data-intensive services, the authors also proposed a database placement strategy based on throughput maximization and two effective database replacement strategies. On the other hand, Huynh et al.[153] propose an innovative DT solution that combines MEC and URLLC to support Metaverse applications. From the perspective of the 3C integration model, this solution optimizes edge caching policies, task offloading strategies, and the allocation of computing and communication resources to address the latency optimization problem in the Metaverse enabled by DT.
III-B3 Summary
The second part discusses the challenges and solutions for resource allocation strategies in the Industrial Metaverse. First, in terms of computing resource allocation, distributed technologies such as cloud computing, MEC, and FL have been introduced to optimize resource distribution due to the varying performance of applications and devices in the Industrial Metaverse. While cloud computing offers powerful computing capabilities, it struggles to meet the low-latency requirements of industrial production; MEC, on the other hand, deploys computing nodes closer to terminal devices, making it suitable for high-concurrency scenarios, although dynamic adjustments to resource allocation strategies are still required to adapt to changing demands. In addition, we explored incentive mechanisms and resource utilization optimization to improve computing resource efficiency. Secondly, in the context of storage resource allocation, dynamic storage strategies and edge storage technologies have been employed to reduce data transmission latency and duplicate processing, particularly for the diverse applications and heterogeneous devices in the Metaverse. However, inconsistency in the sharing of storage resources among edge servers remains an issue. Overall, resource allocation in the Industrial Metaverse must comprehensively consider both computing and storage needs, leveraging distributed technologies, edge computing, and dynamic strategies to enhance platform flexibility and sustainability.
@@ III-C Interoperability
In the context of Industry 5.0, the role of humans in the Industrial Metaverse has become increasingly important. Unlike Industry 4.0, which primarily focuses on automation and machine-driven processes, Industry 5.0 emphasizes human-machine collaboration. This collaboration not only involves interaction between humans and machines at the hardware level but also extends to cognitive cooperation. Human involvement brings unique decision-making abilities, creativity, and adaptability to the process[154]. Moreover, human-AI collaboration allows humans to focus on higher-level tasks, while AI handles repetitive or data-intensive tasks, thereby improving the efficiency and harmony of workflows. At the same time, the Metaverse platform enables human participation by fostering cultural integration and facilitating trustworthy, efficient communication, with interoperability serving as the medium to achieve all of this. As a complex system, the Industrial Metaverse platform integrates various sensors and control subsystems with different hardware architectures and control components. If there is no support for interoperability, user identities and industrial equipment data will be limited to specific VSPs, contradicting the original intention of an interconnected Metaverse[155]. Therefore, interoperability of subsystems within the platform and interoperability across platforms are important considerations. Currently, addressing the interoperability issues in the Industrial Metaverse primarily focuses on three aspects: identity and credential authentication, HCI technology, and interface and integration frameworks.
@@
Identity and credentials define the unique encoding of users and production equipment, while binding identifiers to attributes such as device status and personnel behavior, serving as a foundational element for achieving interoperability. This forms the foundation for achieving interoperability. These attributes can be provided by physical devices in factories or created by enterprise personnel themselves. In different platforms, the same identity and identification correspond to the same static attributes. Gadekallu et al.[156] suggest that BC can facilitate the exchange of data between different sub-Metaverses. They proposed using cross-chain protocols to exchange data on multiple BCs located in different virtual worlds. Therefore, Li et al.[157] propose a MetaOpera, a BC-based interoperability protocol for the Metaverse. It supports cross-chain technology and on-chain/off-chain techniques for the interaction between MetaOpera and decentralized platforms as well as top centralized platforms. Compared to authentication technologies based on Sidechains[158], this method achieves lower interaction latency. Additionally, Ghirmai et al.[159] propose using Self-Sovereign Identity (SSI) technology, which replaces traditional identity verification with distributed identity verification using zero-knowledge proofs. For example, two workers docked on a Metaverse platform can prove each other’s public keys according to the principles of SSI and encrypt their interaction records using pre-shared keys. Each worker holds a unique NFT as a unique identifier. However, the authors also pointed out that the issue of BC storage should not be ignored because controlling and managing data and identities may require different storage mechanisms[160]. Additionally, their scenario only applies to static attributes in the Metaverse and does not involve behavioral objects.
HCI technology, based on identity-authenticated operations, provides users with the ability to interact with virtual environments. Multimodal interaction combines various interaction methods, including motion tracking, haptic feedback, vision, and voice, offering Industrial Metaverse users a natural and intuitive interaction experience for executing complex tasks and operations[161]. Currently, most motion tracking technologies, such as Vicon[162] or Xsens[163], rely on optical markers or inertial measurement units (IMUs). However, these systems are costly to set up and require lengthy calibration times, which smaller companies may struggle to support. To address this, Ponton et al.[164] propose a data-driven method called SparsePoser for body posture reconstruction with sparse data. This method uses six six-degree-of-freedom tracking devices to create avatars. SparsePoser encodes information obtained from sensors along with the user’s static representation and decodes it into full-body postures by reconstructing all joints between end-effectors. Jang et al.[165], on the other hand, focus on reducing equipment costs and proposed a new feature encoder that learns the correlation between historical 3D point cloud data and global and local posture features, achieving real-time motion capture with a single LiDAR. Modeling muscle activity by simultaneously detecting muscle deformation and bioelectrical signals typically requires multiple sensors or even large detection devices, which significantly reduces the wearer’s comfort. To address this, Wang et al.[166] introduce a fully integrated dual-mode wearable system, laying the foundation for a lightweight, dual-modal muscle sensing frontend. Suo et al.[167] present a non-invasive muscle-sensing wearable device that, when combined with machine learning, enables motion-control-based interaction between the physical body and Metaverse avatars.
Another important topic in HCI is haptic technology, as the initial goal of HCI is to enable users to perceive and interact with virtual environments in the same way they would with physical environments. Mel Slater et al.[168] identify two factors that influence this sense of realism: place illusion and plausibility illusion. The former is limited by the capabilities of the interaction devices themselves, while the latter depends on the content of the depicted scene. Due to the limitations in bandwidth and dynamic range, achieving physical stimulation that exactly matches human perception is often impractical. Therefore, Lee et al.[169] suggest using optical or acoustic signals to algorithmically generate vibrations. They categorized the motion into slow and fast, as well as discrete and continuous. Slow motion is relatively easy to reconstruct, but other content often requires a deeper understanding of scene semantics. Kim et al.[170] provide experimental evidence to support this approach. Similarly, research on extracting vibrations from acoustic signals still requires experts for measurement or improvements in the output of suboptimal reconstruction algorithms.
In addition to motion tracking and haptic technology, HCI technology in the Industrial Metaverse also involves digital human simulation. In industrial settings, engineers often use human simulation systems, such as Jack, to assess posture, reachability, and other factors[171]. However, these systems typically focus on body structure and range of motion, overlooking the user’s psychophysiological state. To address this, Eyam et al.[172] introduce the concept of meta-states—digital representations of psychophysiological states—in industrial simulations to reflect the psychophysiological state of digital humans. This allows subjects immersed in virtual reality and professionals operating desktop-based applications to understand the state of human workers. However, due to the high individual variability in psychophysiological states, achieving this simulation poses challenges in terms of algorithmic complexity and computational resource consumption.
TABLE VI: Research on Existing issues in the Industrial Metaverse.
@@
Security & Privacy Sensor-side Industrial Data Source [108] A data misuse prevention mechanism called DAPter. DAPter can substantially raise the bar of the data abuse difficulty with little impact on QoS and overhead. 2021
@@
Inherent Security Vulnerabilities [133] AcoPalm, a palm print recognition system. AcoPalm is resistant to replay and mimicry attacks, and can achieve 96.22% authentication accuracy. 2022
@@
Additionally, during the transition from Industry 4.0 to Industry 5.0, industrial robotics technology faces the need for intelligent and highly integrated development. For industrial robots, Ma et al.[175] propose application scenarios for machine-readable standards within the Metaverse, including resource retrieval, knowledge Q&A, personalized knowledge push, and virtual digital humans. You et al.[176] explore the implementation of virtual-real integration and human-machine collaboration within the Industrial Metaverse. They proposed a design framework for a virtual-real interaction system based on Robot Operating System (ROS) and web technologies, including the design and implementation of a general communication mechanism using ROS, enabling connectivity and real-time data communication between physical robots and virtual models through ROS topic subscriptions. Furthermore, they utilized URDF model transformation technology for model invocation and display.
Finally, as the key to system integration, interfaces and integration frameworks facilitate communication and functional collaboration between different subsystems and platforms. Split learning, as a distributed architecture, involves multiple devices and servers in an industrial platform[177]. It can split the backbone network into multiple sub-networks, execute multiple sub-tasks in parallel on different devices or servers, and then integrate the results. Inspired by this, Rawal et al.[178] introduce a Split-Protocol that provides interoperability services. They assume that critical elements and functions are provided by a meta-server, and the system consists of ten identical meta-servers, each executing specific functions of the Metaverse. Role transitions can be achieved through the Split-Protocol. This protocol with role-switching capability can improve the availability of the Metaverse system. Furthermore, Hashash et al.[174] emphasize the coordination between DTs and Edge Metaverse. A Physical Twin (PT) system running in a large-scale sensing area is replicated as a Cyber Twin (CT) system on MEC servers. The area is divided into distributed subgrids and transmitted remotely to MEC servers. Their designed interactive algorithm utilizes optimal transport theory, considering both computational power and data synchronization issues.
Summary: As a complex system, the Industrial Metaverse needs to address interoperability issues both within subsystems and across platforms. Identity and identification authentication serve as the foundation for interoperability by binding attributes such as device status and human behavior, ensuring consistency across different platforms. BC technology, cross-chain protocols, and distributed identity authentication provide technical support for data exchange and verification. Additionally, distributed architectures such as the Split Protocol, along with the coordination of DTs and edge computing, enable efficient collaboration and data synchronization within the Industrial Metaverse. However, the Industrial Metaverse platforms currently lack a unified standard development model, making it crucial to enhance interoperability and compatibility. Furthermore, only through an open collaborative ecosystem among different companies and industries can the content and value of the Industrial Metaverse be enriched, thereby improving its overall operational experience[179].
@@ IV Standardization
TABLE VII: Comparison of 3D Data Formats in the Metaverse
Format Developer Use Cases Efficiency Open Source Suitability for Industrial
VRML Web3D Consortium Early VR and Web Low Yes Low
glTF 2.0 Khronos Group Metaverse, games, Web High Yes Medium to High (good for visualization and lightweight assets)
USD Pixar Large-scale 3D content creation Medium Yes High (scalable for complex simulations and large datasets)
FBX Autodesk 3D animation, game Low No Medium (useful for detailed models but large file size can be a drawback)
OBJ Wavefront Technologies 3D model exchange Low Yes Low (limited support for complex models or interactions)
COLLADA Khronos Group Cross-tool scene exchange High Yes Medium (suitable for cross-platform exchanges)
@@
Since the concept of the Metaverse was proposed, relevant standardization work has been continuously developed to promote its industrial deployment, ensure interoperability and compatibility between different Metaverse platforms, and reduce the cost of industrial system development and maintenance. While most existing Metaverse standards focus on commercial applications, general characteristics such as interoperability and compatibility can be adapted and expanded to meet the needs of the Industrial Metaverse, providing valuable references and guidance for its development.
From a data perspective, the focus of standardization lies primarily on platform data format standardization and the standardization of data formats between physical and virtual worlds[180]. For instance, interoperability between industrial IoT, industrial cloud platforms, and other systems often requires a common data protocol. Standardized data formats facilitate accurate data transmission, reduce technical barriers for cross-system interactions, and enable developers to create cross-platform tools more efficiently. Since 2008, the MPEG-V standard has defined data exchanges at the interface between the real and virtual worlds to standardize 4D sensory and virtual environments[181]. This standard, comprising seven parts, starts with 23005.1, which outlines an architecture and specifies information representations for interoperability. Parts 23005.2-3 cover syntax and semantics related to sensory device and sensor functions, while parts 23005.5-6 focus on data formats and data types—essential for multimodal data scenarios in the industrial field[182]. In 2017, the Metaverse Standards Forum (MSF) recognized glTF (GL Transmission Format) 2.0 as a benchmark for 3D asset interoperability[183], facilitating the creation and storage of high-fidelity 3D models of industrial equipment or products, enabling efficient cross-platform and multi-context transmission and display. Subsequent data format standards like USD (Universal Scene Description) and COLLADA (Collaborative Design Activity) have further advanced cross-software data interaction and file transfer. Table VII provides a concise comparison of the characteristics of these Metaverse data formats and their applicability in industrial scenarios. In 2019, the IEEE Standards Association launched the IEEE 2888 project, which enhances MPEG-V by standardizing data formats and APIs to synchronize the virtual and physical worlds, improving sensor data collection and actuator control. IEEE 2888 comprises five parts, including specifications for sensor interfaces, actuator interface standards, digital synchronization orchestration between networks and the physical world, VR disaster response system architecture, and evaluation methods. Sections 2888.1-3 provide general technologies for DTs or the Metaverse, while section 2888.4 offers specific applications, such as industrial examples like chemical incident response and fire-fighting training[184]. Its layered system design integrates VR content, motion tracking, sensory feedback, and real-time data processing to deliver comprehensive and immersive emergency response training environments.
From a technical perspective, ISO/IEC/IEEE 24765:2017 defines Metaverse interoperability as the ability of systems and components to exchange and use information[185]. This includes seamless collaboration between agents, system-level data transfer, and object communication and exchange of actionable information. In June 2022, the Khronos Group initiated and established the MSF, bringing together industry giants such as Meta, Microsoft, NVIDIA, and others across sectors like chip manufacturing and gaming. The Forum has amassed over 2,400 companies and standardization organizations, aiming to foster open Metaverse development, interoperability testing, and the deployment of open-source tools. In July 2022, the ISO/IEC JTC 1/SC 29 MPEG WG2 began collecting Metaverse case studies and conducting technical analyses to identify coding and system standards required to support future Metaverse experiences. In January 2023, the IEC Standardization Evaluation Group 15 on Metaverse (SEG 15) was established to formulate a Metaverse standardization roadmap, establish connections with related organizations, and propose further work recommendations. Recently, the IEEE P2048.101 series standards have been developed to provide technical frameworks, components, and main business processes for AR systems on mobile devices in industrial settings[186]. Fig 10 illustrates the logical support relationships among key standards in the standardization of the Industrial Metaverse.
Figure 10: The logical support relationships among key standards in the standardization of the Industrial Metaverse. MPEG-V (2008) laid the foundation for interfacing real and virtual worlds and was later extended by IEEE 2888 (2019) to enhance multimodal synchronization capabilities. This was further supported by IEEE 24765:2017, which ensures consistent terminology across systems. The Metaverse Standards Forum (2017) emphasizes 3D asset interoperability and supports the following data formats: COLLADA (developed by Khronos Group, enabling multi-software compatibility), glTF 2.0 (also provided by Khronos, facilitating efficient and immersive content transmission), and USD (created by Pixar Animation Studios, enabling high-fidelity scene descriptions and cross-platform interactions). Additionally, IEEE P2048.101 offers a framework for VR/AR systems, guiding the integration of immersive technologies.
Overall, the development of Industrial Metaverse standards is still at an early stage. One early example was IBM and Second Life’s effort in 2008, which established a framework for virtual world interoperability using the Open Grid protocol[187]. However, the pressing issue for interoperability extends beyond connecting virtual worlds to encompass the evolution of the entire network, as it transitions to three-dimensional interfaces. While Abilkaiyrkyzy et al.[188] have conducted preliminary studies on small components within industrial platforms, comprehensive and systematic research remains a pressing need.
Given that industrial scenarios involve integrating various sensors and industrial control devices, existing interoperability and compatibility standards need further expansion and refinement, with higher flexibility and modular design introduced to meet diverse industrial scenarios and industry needs. As the Industrial Metaverse gradually materializes, its accompanying standards in interoperability, security and privacy, openness, scalability, and cross-industry collaboration will become increasingly robust, forming a comprehensive standard system.
@@ V-A Flexible Communication Network Architecture
The seamless internet connectivity of the Metaverse largely relies on reliable broadband services, with holographic communication requiring real-time operation, sufficient bandwidth, and throughput[189]. While low latency and high reliability are the primary advantages of 5G/6G services, consistent latency requires a flexible end-to-end architecture. This is because the Industrial Metaverse requires real-time communication of video, audio, and data. Currently, there is limited research on the network architecture and signal processing aspects of 6G in the context of the Metaverse. This calls for a strong foundation in wireless communication theory, signal processing, and expertise in network architecture design.